if a boy si 24 years younger than him father after 2 years their sum of ages will be 40,what is the age of father?

Answer by mananth(16949)

let the father's age be x now
son's age = x-24
.
After 2 years
father will be x+2
son will be x-24 +2
Their sum = 40
x+2 +x-24+2 = 40
2x-20=40
2x = 60
x= 30
Father's age = 30 years.